# Zebrafix: Mitigating Memory-Centric Side-Channel Leakage via Interleaving

### v1

@@ 5.2 Adjusting Compiler Intrinsics
Examples of memory-altering intrinsics are $\mathtt{memcpy}$ and $\mathtt{memset}$, which we replace with own implementations in Zebrafix. As the 16-byte blocks that are built with Zebrafix contain varying amounts of data (e.g., 1 byte for chars, 4 byte for short integers, 8 byte for long integers), it does not suffice to adjust the amount of bytes copied or set by a fixed factor. Instead, the correct number of elements of a specific primitive type to be copied has to be determined. With that, the data can properly be extracted such that a suitable interleaved struct for the copy or set destination can be generated. The struct is then used for inserting new instrumented writes to memory. In addition, the counter values must not only be copied, but also incremented to prevent leakage from multiple write accesses with the same counter value.
5.3 Instrumentation Aspects for Interleaving
@@
First, a global counter has to be added so that we can ensure that in all contexts of the program, freshness is integrated into a memory write. Else, if for example a function accesses the same address in separate calls, it could reuse the same old counter value. The counter value gets initialized with a random value and is then incremented on each memory write to enforce ciphertext changes in each written block. Similar to the counter, as a basic part of the mitigation, we also need to add all interleaved structs that we need later on to insert proper handling for each primitive data type (e.g., integers with varying bit width, see Figure 7).
@@
Functions to be instrumented are determined by propagating a $\mathtt{clang}$ function attribute of a base function along the call graph of that function. Afterward, the call to the base function is updated to now call its instrumented equivalent that only has instrumented child functions. Inside those instrumented functions, all address computations (called $\mathtt{GEP}$s in LLVM), memory loads and stores (alongside instructions that contain implicit loads and stores like constant expression $\mathtt{GEP}$s) and allocations have to be instrumented to follow the layout of interleaved and thereby extended memory. Global variables and named structs also have to be adjusted to an interleaved format.
@@
A share of the memory being extended by adding interleaving are heap allocations. In order to ensure enough available space for adding interleaved structs on the heap, we extend the memory allocated by calls to $\mathtt{malloc}$ and related allocation functions. As $\mathtt{malloc}$ does not use type information, we assume the “worst” case interleaved struct that contains one byte of data in each 16-byte block. We do not add the same SVF pointer analysis pass we added for $\mathtt{memcpy}$ or $\mathtt{memset}$ to lower the amount of manual intervention needed during compilation for fixing unknown cases. While compiler intrinsics for copying or setting values can break with a default handling, the handling of heap allocations works stable when setting the default to the byte-wise handling.
5.4 Binary Level Verification
@@
To ensure that the LLVM middle end level instrumentation is not removed by compiler passes that occur later in the compilation process, and that inserted stack spills do not leak any secret-dependent values, we implemented a binary level verification tool. The verification tool is a Pintool [27] that traces all memory writes of the instrumented binary and manages a set of all written values per address. Whenever there is a repetition of written values within a set, the corresponding instruction offsets and written values are logged to a trace file so that the developer can manually check them for potential leakages.
6 Zebrafix Evaluation
@@
In order to categorize the presented approach of Zebrafix as a mitigation against ciphertext side-channel leakage, we assess the security and performance in comparison to the Cipherfix masking approach.
@@
We perform our evaluation on an Intel 4th Gen Xeon SP Server (Sapphire Rapids). We base our evaluation on the reference implementations of primitives in $\mathtt{libsodium}$-1.0.20 and $\mathtt{mbedTLS}$-3.6.0 with conformity to Section 9.2, i.e., we preempt target-specific primitive versions which include inline assembly that is out-of-scope for compiler level analysis. We disabled vectorization flags for building the merged bitcode file that is instrumented by the Zebrafix LLVM passes.
6.1 Compile Time and Memory Consumption
@@
The compile time is composed of different compilation steps. We first build a non-instrumented version of the library composed of various bitcode files. As this step is the same for both non-instrumented and instrumented versions, we exclude it from the measurements. We then measure the time it takes to combine the necessary bitcode files into a non-instrumented version as well as into an instrumented version. The compile time that is needed for building instrumented binaries using a pre-built library runs up to eight seconds while the sequential non-instrumented build process takes around three seconds.
@@
Table I: Overhead in terms of compilation time, code size and memory consumption. Compilation time and code size are related to the library version containing certain instrumented primitives, whereas memory usage is estimated for the execution of the primitive in isolation.
@@
6.2 Security & Runtime Overhead
@@
As Zebrafix provides nonce-based randomization within the selected block size, the resulting encryption is fully probabilistic (analogous to CTR-mode encryption). With a block size of 16 bytes, of which 8 byte are counters, a repetition occurs only after $2^{64}$ write accesses. In our implementation, every memory write is protected. The exception to this are stack spills that are introduced after the LLVM IR instrumentation because of too much register pressure in the vector registers for storing intermediate variables. The stack spills are displayed to the developer during the compile process and can further be checked at the binary level (cf. Section 5.4). That way, the developer can determine whether secret data would be spilled to the stack and can apply some source level mitigations like storing the affected data as global variables instead.
@@
The primitives for the evaluation were chosen according to the following criteria: $\mathtt{EdDSA}$ has been shown to be vulnerable against ciphertext side-channel attacks [25, 42]. Apart from that, we include $\mathtt{Base64}$ as it was the fastest target in prior ciphertext side-channel mitigation evaluations, and $\mathtt{SHA512}$ as another fast target for comparison. We depict the $\mathtt{Base64}$ decoding in Table II; the results for the encoding are similar. All targets include zeroing of the resp. memory when freeing resources to avoid leakage of data through memory reuse. As mentioned in Section 9.2, we exclude targets that contain inline assembly or features that the current proof-of-concept does not support, like $\mathtt{AES}$ vector instructions. In the latter case, the computation mainly happens in registers, so leaking spots are rare and can be subject to manual adjustments.
@@
We measured the runtime overhead of Zebrafix for each primitive in relation to an execution of the baseline version that is using our LLVM with vector register-style stack spilling. The baseline version and the instrumented version likewise share the same build parameters. In line with prior work, we thereby estimate the overhead that our instrumentation adds rather than the total execution time. We add ten rounds of warm up and include 1000 rounds of measuring each target’s execution overhead. Furthermore, we repeat the aforesaid procedure for ten times and average all runs for more precise estimations of the overhead. The Zebrafix instrumentation adds an average overhead factor of 1.2×1.2$\times$1.2 × over all targets, whereas the most balanced version of Cipherfix adds an average overhead of 4.2×4.2$\times$4.2 × and the most secure version an overhead of 21.8×21.8$\times$21.8 ×. The detailed results are shown in Table II.
@@
Although the memory tracking Pintool described in Section 5.4 does not provide information on the exploitability of potential leakage, code can be divided into known vulnerable functions where developers need to pay more attention during manual analyses, and other functions containing colliding written values. Filtering the instruction offsets has verified that the known vulnerable functions do not contain any leaky memory writes. Other leakages that need to be checked in other functions are $\mathtt{push}$ instructions and about 25 unprotected memory writes in the described targets that are not interleaved 16-byte writes.
6.3 Comparison to Cipherfix Masking Approach
@@
Even for the fastest primitive with the weakest security level in the Cipherfix evaluation, $\mathtt{Base64}$ from $\mathtt{mbedTLS}$ and Cipherfix-Fast with $\mathtt{AES}$, Zebrafix reaches a similar overhead factor. In Cipherfix, only secret-dependent instructions are protected. Thus, Zebrafix’s performance impact could be reduced even further by adding secret tracking and other optimizations, e.g., allowing the general use of vector instructions in combination with the interleaved memory writes.
In general, our results suggest that our proposed interleaving approach is more efficient than the data masking approach taken by Cipherfix. This is due to the fact that randomness generation introduces a significant overhead that we can avoid. The remaining speedup of Zebrafix over Cipherfix can be explained with a more optimized binary that was instrumented on compiler level, instead of relying on binary rewriting. From this, we conclude that implementing an interleaving-based ciphertext side-channel mitigation in the compiler middle end can enhance performance and security compared to masking-based binary-level approaches.
Observation 1 To balance performance and security, a mitigation in the compiler middle end can take advantage of memory layout customization and ease portability across architectures.
@@
$\mathtt{Ed25519}$ $39.1$$\times$ $5.5$$\times$ $8.6$$\times$ $1.3$$\times$
$\mathtt{SHA512}$ $21.6$$\times$ $1.6$$\times$ $2.4$$\times$ $1.1$$\times$
@@
$\mathtt{Base64}$ $4.6$$\times$ $1.2$$\times$ $1.6$$\times$ $1.2$$\times$
7 Preventing Leakages from Silent Stores
With the discussion of ciphertext side-channel defenses in mind, we now elaborate on interconnections to silent stores to show how these memory-centric leakages can be generalized to provide more holistic defenses. Per se, all attacks described in Section 2 are based on different mechanisms to exploit leakage. However, despite originating from different sources, all of these leakages can be considered memory-centric side-channels and thus be thwarted by defenses protecting against the whole class: If the data was not architecturally written to memory in a leaky way in the first place, the different leakage mechanisms are no longer effective. In this section, we outline the interconnection of ciphertext side-channel and silent store leakages and compare the effectiveness of Zebrafix with silent store mitigations.
@@ 7 Preventing Leakages from Silent Stores
In the ciphertext side-channel leakage scenario, an attacker can observe deterministic patterns of ciphertext changes of victim VM data and thereby derive information about secret keys. The attacker model of ciphertext side-channel leakage is the strongest one to be considered. In the other memory-centric leakage scenarios, we assume that microarchitectural optimizations that are currently on the rise, like the ones discussed in Section 2, will gain more importance and see implementation in common processors. We therefore assume a strong attacker to err on the side of caution. The attacker is able to bring the system into a state where the best conditions for maximal leakage through microarchitectural optimizations have been established; thus we assume the worst case leakage in each channel to be possible. [^4] At the time of writing, we are not aware of systems that already feature the microarchitectural optimizations in question, apart from the ones that are linked in Section 2. The assumptions on code that is instrumented are in line with the ones presented in Section 3.1.
7.1 Interconnecting Silent Store and Ciphertext Side-Channel Leakage
@@
In CipherH [15], the authors propose a way to detect ciphertext side-channel leakage with the help of symbolic execution. Consecutive secret-dependent memory writes are checked for whether the values to be written are always the same, always different or possibly different. Thereby, ciphertext side-channel leakage is found for the case of two consecutive memory writes of the same secret-dependent variable to the same address if the value of the variable may vary. However, the leakage model does not cover all leakages through the ciphertext side-channel. Missing cases include interleaved memory writes of secret values to the same address or leakage through observations of ciphertext changes between a known public value and a secret value (e.g., replacing an all-zero public value with a secret that can either be zero as well or some other value). Thus, CipherH only finds a case of silent store leakage, where two consecutive secret-dependent memory writes with possibly different values occur.
@@
The silent store mitigation cio [19] adds a memory write with a provably different value between two consecutive writes. Thereby, the leakage through silent stores is thwarted. The proposed mitigation still allows observing the ciphertext patterns that lead to ciphertext side-channel leakage as the attacker just has to monitor more memory writes and remove the dummy writes (the ones that hinder silent stores) in between. Even worse, the incomplete modeling of ciphertext side-channel leakage in CipherH would incorrectly state that code instrumented with a cio-style approach does not leak in the ciphertext side-channel scenario.
@@
However, any defense against ciphertext side-channel leakage also protects against leakage from silent stores if the option of silencing the store is checked for the same chunk width as the inserted memory write width. In the case of rotating the location a data element is written to, and under the assumption that the memory is initialized properly, the memory addresses are not reused for the same data and are thereby protected against leakage from silent stores. In the case of a probabilistic software encryption layer through masking or interleaving, all written values are different due to the probabilistic encryption, and there are no stores that can deterministically be suppressed by microarchitectural optimizations.
Observation 2 Adding additional freshness to data before writing it to memory prevents leakage through both ciphertext side-channels and silent stores.
7.2 Comparison of Zebrafix to Silent Store Mitigations

### v2

@@ V-B Adjusting Compiler Intrinsics
Examples of memory-altering intrinsics are $\mathtt{memcpy}$ and $\mathtt{memset}$, which we replace with own implementations in Zebrafix. As the 16-byte blocks that are built with Zebrafix contain varying amounts of data (e.g., 1 byte for chars, 4 byte for short integers, 8 byte for long integers), it does not suffice to adjust the amount of bytes copied or set by a fixed factor. Instead, the correct number of elements of a specific primitive type to be copied has to be determined. In the Zebrafix implementation, we use SVF to infer the types of (source and) destination objects for the memory-altering intrinsics. With that, the data can properly be extracted such that a suitable interleaved struct for the copy or set destination can be generated. The struct is then used for inserting new instrumented writes to memory. In addition, the counter values must not only be copied, but also incremented to prevent leakage from multiple write accesses with the same counter value. Thereby, even with multiple copies of the buffer, the counters are always incremented to avoid repeating ciphertexts.
V-C Instrumentation Aspects for Interleaving
@@
First, for having access to a global counter in all contexts of the program, a vector register reserved for the counter is initialized with a random counter value generated by $\mathtt{rdrand}$. For easing the handling of global variables, we also add an initializer counter with a value that is known at compile time. This initializer gets replaced with the vectorized counter during the first memory write to the global variable. Similar to the counter, as a basic part of the mitigation, we also need to add all needed interleaved structs to insert proper handling for each primitive data type (e.g., integers with varying bit width, see Figure 6).
@@
Functions to be instrumented are determined by propagating a $\mathtt{clang}$ function attribute of a base function along the call graph of that function. Afterward, the call to the base function is updated to now call its instrumented equivalent that only has instrumented child functions. Inside those instrumented functions, all address computations (called $\mathtt{GEP}$s in LLVM), memory loads and stores (alongside instructions that contain implicit loads and stores like constant expression $\mathtt{GEP}$s) and allocations have to be instrumented to follow the layout of interleaved and thereby extended memory. Global variables and named structs also have to be adjusted to an interleaved format. An example of the instrumentation of a memory write including the counter update can be found in Figure 7.
Figure 7: Instrumentation of a 1-byte store instruction. The value to be written is combined with a counter via register-only operations. Finally, a single store instruction writes both value and counter.
@@
A share of the memory being extended by adding interleaving are heap allocations. In order to ensure enough available space for adding interleaved structs on the heap, we extend the memory allocated by calls to $\mathtt{malloc}$ and related allocation functions. As $\mathtt{malloc}$ does not use type information, we assume the “worst” case interleaved struct that contains one byte of data in each 16-byte block. An example of an instrumented version of $\mathtt{malloc}$ is given in Figure 8. We do not add the same SVF pointer analysis pass for type inference we added for $\mathtt{memcpy}$ or $\mathtt{memset}$ to lower the amount of manual intervention needed during compilation for fixing unknown cases. If type inference is needed to reduce memory overhead for certain examples, this could be adjusted. While compiler intrinsics for copying or setting values can break with a default handling, the handling of heap allocations works stable when setting the default to the byte-wise handling. Combined with the extension of the allocated heap space, the writes to heap objects are instrumented to not only include the writes of original data but also the 16-byte writes that combine original data and an incremented counter.
Figure 8: Instrumentation of a $\mathtt{malloc}$ call that allocates a byte array. The allocation size is adjusted by the compiler to fit the additional counters and padding.
V-D Binary-Level Verification
@@
To ensure that the LLVM middle end level instrumentation is not removed by compiler passes that occur later in the compilation process, and that inserted stack spills do not leak any patterns of repeating values, we implement a binary-level verification tool. The verification tool is a Pintool [23] that traces all memory writes of the instrumented binary and manages a set of all written values per address. Whenever there is a repetition of written values within a set, the corresponding instruction offsets and written values are logged to a trace file so that the developer can manually check them for potential leakages in accordance with known leaking algorithms. That way, the developer can determine whether secret data would be spilled to the stack and can apply some source-level mitigations like storing the affected data as global variables instead. Afterward, the compilation and binary-level check is repeated until a non-leaky version is produced (cf. Figure 5).
VI Zebrafix Evaluation
@@
In order to categorize the presented approach of Zebrafix as a mitigation against ciphertext side-channel leakage, we assess the security and performance in comparison to the Cipherfix masking approach. We perform our evaluation on a system with an Intel 4th Gen Xeon Gold 6438Y+ CPU (Sapphire Rapids) with $512$ GiB of DDR5 and local SSD storage. We base our evaluation on target primitives from $\mathtt{libsodium}$-1.0.20 and $\mathtt{mbedTLS}$-3.6.0.
VI-A Compile Time and Memory Consumption
@@
To estimate the additional memory consumption and the additional time required to run the instrumentation process, we base our evaluation on the reference implementations of primitives in $\mathtt{libsodium}$ and $\mathtt{mbedTLS}$. The reference implementations conform to Section VIII, i.e., we preempt target-specific primitive versions that include inline assembly, which is out of scope for compiler-level analysis. We have disabled vectorization flags for building the merged bitcode file that is instrumented by the Zebrafix LLVM passes. While a production-level version of the code may have less overhead in code size due to optimized inline assembly and less compilation time due to parallel build processes, we can better estimate the additional overhead. The library build process happens only once, and we verify that the order of magnitude of the build remains the same. To measure the runtime overhead (which is present within each execution), we compare to a vectorized baseline.
@@
The compile time is composed of different compilation steps. For measurements of compile time and memory consumption, we first build a non-instrumented version without vectorization of the library composed of various bitcode files. We measure the time it takes to combine the necessary bitcode files into a non-instrumented version as well as into an instrumented version. The compile time that is needed for building instrumented binaries using a pre-built library runs up to eight seconds while the sequential non-instrumented build process takes around three seconds.
@@
Table I: Overhead in terms of compilation time, code size and memory consumption. Compilation time and code size are related to the library version containing certain instrumented primitives (baseline given in the library rows), whereas memory usage is estimated for the execution of the primitive in isolation.
@@
VI-B Runtime Overhead and Security
@@
As Zebrafix provides nonce-based randomization within the selected block size, the resulting encryption is fully probabilistic (analogous to CTR-mode encryption). With a block size of 16 bytes, of which 8 byte are counters, a repetition occurs only after $2^{64}$ write accesses. In our implementation, every memory write is protected. The exception to this are stack spills that are introduced after the LLVM IR instrumentation because of too much register pressure in the vector registers for storing intermediate variables. The stack spills are displayed to the developer during the compile process and can further be checked at the binary-level (cf. Section V-D).
@@
The primitives for the evaluation were chosen according to the following criteria: $\mathtt{EdDSA}$ has been shown to be vulnerable against ciphertext side-channel attacks [21, 37]. Apart from that, we include $\mathtt{Base64}$ as it was the fastest target in prior ciphertext side-channel mitigation evaluations, and $\mathtt{SHA512}$ as another fast target for comparison. We depict the $\mathtt{Base64}$ decoding in Table II; the results for the encoding are similar. All targets include zeroing of the resp. memory when freeing resources to avoid leakage of data through memory reuse. As mentioned in Section VIII, we exclude target versions that contain inline assembly or features that the current proof-of-concept does not support, like $\mathtt{AES}$ vector instructions. In the latter case, the computation mainly happens in registers, so leaking spots are rare and can be subject to manual adjustments.
@@
We measured the runtime overhead of Zebrafix for each primitive in relation to an execution of the baseline version that is using a plain LLVM version. The baseline version and the instrumented version share the same build parameters where possible (standard parameters with additional –disable-asm) and include disabling the vectorization for the instrumented version to ensure that all instructions can be properly instrumented by Zebrafix. The Zebrafix instrumentation adds an average overhead factor of 1.3×1.3$\times$1.3 × over all targets, whereas the most balanced version of Cipherfix adds an average overhead of 4.2×4.2$\times$4.2 × and the most secure version an overhead of 21.8×21.8$\times$21.8 ×. The detailed results are shown in Table II.
@@
Although the memory tracking Pintool described in Section V-D does not provide information on the exploitability of potential leakage, code can be divided into known vulnerable functions where developers need to pay more attention during manual analyses, and other functions containing colliding written values. Filtering the instruction offsets has verified that the known vulnerable functions do not contain any leaky memory writes. Other leakages that need to be checked in other functions are $\mathtt{push}$ instructions and about 25 unprotected memory writes in the described targets that are not interleaved 16-byte writes.
VI-C Comparison to Cipherfix Masking Approach
@@
Even for the fastest primitive with the weakest security level in the Cipherfix evaluation, $\mathtt{Base64}$ from $\mathtt{mbedTLS}$ and Cipherfix-Fast with $\mathtt{AES}$, Zebrafix reaches a similar overhead. In Cipherfix, only secret-dependent instructions are protected. Thus, Zebrafix’s performance impact could be reduced even further with secret tracking and other optimizations, e.g., allowing the use of vector instructions in combination with interleaved memory writes.
In general, our results suggest that our proposed interleaving approach is more efficient than the data masking approach taken by Cipherfix. This is due to the fact that randomness generation introduces a significant overhead that we avoid. The remaining speedup of Zebrafix over Cipherfix can be explained with a more optimized compiler-built binary, instead of relying on binary rewriting. From this, we conclude that implementing an interleaving-based ciphertext side-channel mitigation in the compiler middle end can enhance performance and security compared to masking-based binary-level approaches.
@@
$\mathtt{Ed25519}$ $39.1$$\times$ $5.5$$\times$ $8.6$$\times$ $1.6$$\times$
$\mathtt{SHA512}$ $21.6$$\times$ $1.6$$\times$ $2.4$$\times$ $1.3$$\times$
@@
$\mathtt{Base64}$ $4.6$$\times$ $1.2$$\times$ $1.6$$\times$ $1.1$$\times$
VII Preventing Leakages from Silent Stores
With the discussion of ciphertext side-channel defenses in mind, we now explore connections to silent stores to show how these memory-centric leakages can be generalized to more holistic defenses. Per se, all attacks described in Section II are based on different mechanisms to exploit leakage. However, despite originating from different sources, all of these leakages can be considered memory-centric side-channels and thus be thwarted by defenses protecting against the whole class: If the data was not architecturally written to memory in a leaky way in the first place, the different leakage mechanisms are no longer effective. In this section, we outline the interconnection of ciphertext side-channel and silent store leakages and compare the effectiveness of Zebrafix with silent store mitigations.
@@ VII Preventing Leakages from Silent Stores
In the ciphertext side-channel leakage scenario, an attacker can observe deterministic patterns of ciphertext changes of victim VM data and thereby derive information about secret keys. The attacker model of ciphertext side-channel leakage is the strongest one to be considered. In the other memory-centric leakage scenarios, we assume that microarchitectural optimizations that are currently on the rise, like the ones discussed in Section II, will gain more importance and see implementation in common processors. We therefore assume a strong attacker to err on the side of caution. The attacker is able to bring the system into a state where the best conditions for maximal leakage through microarchitectural optimizations have been established; thus we assume the worst case leakage in each channel to be possible. [^4] At the time of writing, we are not aware of systems that already feature the microarchitectural optimizations in question, apart from the ones that are linked in Section II. The assumptions on code that is instrumented are in line with the ones presented in Section III-A.
VII-A Interconnecting Silent Stores and Ciphertext Side-Channel Leakage
@@
In CipherH [12], the authors propose to use symbolic execution to detect ciphertext side-channel leakage. Consecutive secret-dependent memory writes are checked for whether the values to be written are always the same, always different or possibly different. Thereby, ciphertext side-channel leakage is found for the case of two consecutive memory writes of the same secret-dependent variable to the same address if the value of the variable may vary. However, the leakage model does not cover all ciphertext side-channel leakages. Missing cases include interleaved memory writes of secret values to the same address or leakage through observations of ciphertext changes between a known public value and a secret value (e.g., replacing an all-zero public value with a secret that can either be zero as well or some other value). Thus, CipherH only finds a case of silent store leakage, where two consecutive secret-dependent memory writes with possibly different values occur.
@@
The silent store mitigation cio [15] adds a memory write with a provably different value between two consecutive writes. Thereby, the silent store leakage is thwarted. The proposed mitigation still allows observing the ciphertext patterns that lead to ciphertext side-channel leakage as the attacker just has to monitor more memory writes and remove the dummy writes (the ones that hinder silent stores) in between. Even worse, the incomplete modeling of ciphertext side-channel leakage in CipherH would incorrectly state that code instrumented with a cio-style approach does not leak in the ciphertext side-channel scenario.
@@
However, any ciphertext side-channel defense also protects against silent store leakage if the option of silencing the store is checked for the same chunk width as the inserted memory write width. In the case of rotating the location a data element is written to, and under the assumption that the memory is initialized properly, the memory addresses are not reused for the same data and are thereby protected against leakage from silent stores. In the case of a probabilistic software encryption layer through masking or interleaving, all written values are different due to the probabilistic encryption, and there are no stores that can deterministically be suppressed by microarchitectural optimizations.
VII-B Comparison of Silent Store Mitigations